# Rate-In: Information-Driven Adaptive Dropout Rates for Improved Inference-Time Uncertainty Estimation

### v3

@@ 4.2.1 Synthetic Data Analysis
We first evaluate Rate-In on the synthetic regression task to validate our approach under controlled conditions. Figure  3(a) shows the efficiency of uncertainty estimation across increasing noise levels ($\sigma\in\{0.1,0.2,0.3,0.4,0.5\}$) with a fixed training set size ($N=100$). The efficiency is quantified using the ratio of interval width to empirical coverage; lower ratios indicate more precise uncertainty estimates without sacrificing reliability. Rate-In consistently maintains a lower ratio across all noise levels compared to baseline methods.
Figure  3(b) illustrates the impact of training set size under a fixed noise level ($\sigma=0.01$). Rate-In outperforms other methods across all training sample sizes, achieving more efficient uncertainty estimation.The results demonstrate Rate-In’s ability to adapt dropout rates based on local uncertainty characteristics, leading to more efficient and reliable uncertainty estimation compared to static dropout approaches. The method shows particular advantages in scenarios with heteroscedastic noise and limited training data.
@@
Figure 3: Rate-In is more efficient in uncertainty estimation. The ratio of uncertainty interval width to 95% coverage. Lower ratios indicate more efficient uncertainty estimation. (a) Varying noise levels; (b) Varying number of training points at fixed noise level $(\sigma=0.01)$.
4.3 Classification Tasks
Table  1 summarizes the performance metrics for each classification task and dropout policy. Rate-In consistently achieves higher accuracy and better uncertainty calibration compared to baseline methods. Specifically, for the PathMNIST dataset, Rate-In maintains high performance across different levels of information loss, as illustrated in Figure 4. At a 0.05 dropout rate, Rate-In maintains full model accuracy for both TissueMNIST and PathMNIST datasets, while other methods show significant degradation.
@@ 4.3 Classification Tasks
Table 1: Rate-In maintains near-original model accuracy even at high dropout rates, while baseline methods show severe degradation. Classification performance across three medical imaging datasets, reported as Accuracy (ACC) and AUARC. Results demonstrate that Rate-In consistently outperforms conventional dropout strategies, particularly at higher dropout rates. Higher AUARC values (in parentheses) indicate better uncertainty estimation quality.
Dataset Method ACC Full Model Dropout Rate
0.05 0.10 0.20
TissueMNIST Constant .68 .53 (.67) .31 (.37) .09 (.25)
@@
Rate-In .68 (.84) .64 (.80) .48 (.59)
PathMNIST Constant .90 .87 (.96) .75 (.91) .41 (.56)
@@
Rate-In .90 (.98) .88 (.97) .75 (.91)
BloodMNIST Constant .97 .95 (.99) .91 (.98) .58 (.75)
@@
Rate-In .96 (.99) .93 (.99) .82 (.92)
Figure 4: Rate-In maintains high classification accuracy even at aggressive dropout rates where conventional methods degrade. Classification accuracy on the PathMNIST dataset comparing dropout strategies across increasing rates. Traditional methods show sharp decline, while Rate-In preserves performance by adaptively adjusting dropout patterns based on local feature importance.
4.4 Segmentation Tasks
Table 2: Rate-In achieves superior uncertainty estimation across all anatomical zones while preserving or improving segmentation accuracy compared to best benchmarks. Comprehensive evaluation across different anatomical regions in MRI and CT modalities, showing DSC, ECE, and BUC at varying dropout rates. Rate-In demonstrates consistent improvements in uncertainty metrics (lower ECE, higher BUC) while maintaining competitive DSC scores, with particularly strong performance in challenging regions like the transitional zone and tumor boundaries.
Zone Modality Method Full Model DSC (Full Model, % change) ECE ×$\times 10^{-3}$ BUC
DSC p=0.01 p=0.05 p=0.10 p=0.01 p=0.05 p=0.10 p=0.01 p=0.05 p=0.10
@@ 4.4 Segmentation Tasks
Best Benchmark +0.17% -1.90% -1.73% 1.80 1.80 1.90 0.60 0.55 0.53
Rate-In’s advantages are particularly evident in medical image segmentation, where accurate uncertainty estimation is crucial. As illustrated in Figure 1, Rate-In generates uncertainty maps that precisely highlight anatomically challenging regions, especially along tissue boundaries where expert radiologists often face ambiguity. In contrast, constant dropout produces diffuse uncertainty patterns at higher dropout rates, obscuring areas of genuine uncertainty.
@@
  • Error Correlation: Rate-In’s high-uncertainty regions strongly correlate with actual prediction errors, indicating better calibration and enhancing trust in the model’s predictions.
@@
These qualitative observations are supported quantitatively by performance metrics across our test set. Table 2 summarizes the metrics for each segmentation task and dropout policy. Rate-In consistently outperforms benchmarks in uncertainty metrics such as ECE and BUC across most anatomical zones. For instance, in regions with inherent boundary ambiguity, Rate-In shows enhanced capability in quantifying uncertainty at region interfaces, evidenced by higher BUC scores in the Transitional zone. This improved boundary characterization is critical for clinical applications where precise delineation between overlapping anatomical zones impacts diagnostic accuracy.
The results demonstrate that Rate-In’s information-theoretic approach to adjusting dropout rates produces uncertainty estimates that are both more accurate and more clinically meaningful. The method’s ability to maintain focused uncertainty patterns while preserving anatomical detail makes it particularly valuable for medical imaging applications where precision and reliability are paramount.
Figure 5 illustrates the impact of dropout rates on information loss across network layers. In the pre-trained U-Net model, the decoder layers experience a significant decline in information, with a reduction of approximately 40% as dropout rates increase from 5% to 20%. In contrast, the encoder layers are less affected, showing smaller changes in structural similarity with varying dropout rates. This suggests that Rate-In effectively allocates dropout rates to minimize information loss where it is most critical.
Figure 5: Decoder layers are more sensitive to information loss from dropout than encoder layers, demonstrating the need for layer-specific dropout rates. Analysis of information retention across U-Net layers as dropout rates increase. Decoder layers show significant information loss, while encoder layers maintain higher information. This asymmetric behavior validates Rate-In’s layer-adaptive approach, which can allocate lower dropout rates to sensitive decoder layers while maintaining higher rates in more robust encoder layers for efficient uncertainty estimation.

### v4

@@ 4.2.1 Synthetic Data Analysis
We first evaluate Rate-In on the synthetic regression task to validate its performance under controlled conditions. Fig. 3(a) shows the efficiency of uncertainty estimation across increasing noise levels ($\sigma\in\{0.1,0.2,0.3,0.4,0.5\}$) with a fixed training set size ($N=100$). Efficiency is measured as the ratio of interval width to empirical coverage; lower ratios indicate more precise uncertainty estimates without compromising reliability. Rate-In maintains a consistently lower ratio across all noise levels compared to baselines.
Fig. 3(b) shows the impact of training set size under a fixed noise level ($\sigma=0.01$). Rate-In outperforms other methods across all sample sizes, achieving more efficient uncertainty estimation.The results highlight Rate-In’s ability to adapt dropout rates based on local uncertainty characteristics, leading to more efficient and reliable estimates than other dropout baselines. The method is particularly advantageous in scenarios with heteroscedastic noise and limited training data. Sensitivity analysis on training size, noise, and convergence is in Appendix B, Sections  7.1 and  7.5.
@@
Figure 3: Rate-In is more efficient in uncertainty estimation. The ratio of uncertainty interval width to 95% coverage. Lower ratios indicate greater efficiency. (a) Varying noise levels; (b) Varying number of training points at fixed noise level. $(\sigma=0.01)$.
4.2.2 Classification Tasks
Table 1 presents performance metrics for each classification task and dropout policy. Rate-In consistently outperforms baselines in accuracy and uncertainty calibration. Specifically, for the PathMNIST dataset, Rate-In maintains high performance across varying information loss levels. At a 0.05 dropout rate, Rate-In preserves full model accuracy for both TissueMNIST and PathMNIST datasets, while other methods degrade significantly.
@@ 4.2.2 Classification Tasks
Table 1: Rate-In preserves near-original model accuracy even at high dropout rates, while baseline methods degrade significantly. Classification performance across medical imaging tasks is reported as Accuracy (AUARC). Rate-In consistently outperforms traditional dropout methods, particularly at higher dropout rates. Higher AUARC values indicate better uncertainty estimation.
Dataset Method Accuracy Accuracy (AUARC)
Full Model $p\text{,}\epsilon{\text{=}}0.05$ $p\text{,}\epsilon{\text{=}}0.10$ $p\text{,}\epsilon{\text{=}}0.20$
TissueMNIST Rate-In .68 .68 (.84) .64 (.80) .48 (.59)
Constant .53 (.67) .31 (.37) .09 (.25)
@@
PathMNIST Rate-In .90 .90 (.98) .88 (.97) .75 (.91)
Constant .87 (.96) .75 (.91) .41 (.56)
@@
BloodMNIST Rate-In .97 .96 (.99) .93 (.99) .82 (.92)
Constant .95 (.99) .91 (.98) .58 (.75)
@@
4.2.3 Segmentation Tasks
Table 2: Rate-In achieves superior uncertainty estimation across all anatomical zones while preserving or improving segmentation accuracy compared to top benchmarks. Comprehensive MRI and CT evaluations show DSC, ECE, and BUC across varying dropout rates. Rate-In consistently improves uncertainty metrics (lower ECE, higher BUC) while maintaining competitive DSC, excelling in challenging regions like the transitional zone and tumor boundaries. Results for all dropout methods are in Appendix B, Table 4.
Zone Modality Method DSC DSC (Full Model, % change) ECE ×$\times 10^{-3}$ BUC
Full Model $p\text{,}\epsilon{\text{=}}0.01$ $p\text{,}\epsilon{\text{=}}0.05$ $p\text{,}\epsilon{\text{=}}0.10$ $p\text{,}\epsilon{\text{=}}0.01$ $p\text{,}\epsilon{\text{=}}0.05$ $p\text{,}\epsilon{\text{=}}0.10$ $p\text{,}\epsilon{\text{=}}0.01$ $p\text{,}\epsilon{\text{=}}0.05$ $p\text{,}\epsilon{\text{=}}0.10$
@@ 4.2.3 Segmentation Tasks
Benchmark +0.17% -1.90% -1.73% 1.80 1.80 1.90 0.60 0.55 0.53
Rate-In is highly effective in medical image segmentation, where accurate uncertainty estimation is crucial. As shown in Figure 1, Rate-In generates uncertainty maps that accurately highlight anatomically challenging regions, especially along tissue boundaries where radiologists face ambiguity. In contrast, constant dropout diffuses uncertainty at higher dropout rates, obscuring true areas of uncertainty.
@@
  • Error Correlation: Rate-In’s high-uncertainty regions strongly align with prediction errors, indicating better calibration and enhancing trust in the model’s predictions.
@@
These qualitative observations are quantitatively supported by test set performance metrics. Table 2 summarizes results for each segmentation task and dropout policy. Rate-In consistently outperforms benchmarks in uncertainty metrics such as ECE and BUC across most anatomical zones. For instance, in regions with boundary ambiguity, Rate-In enhances uncertainty quantification at region interfaces, as reflected by higher BUC scores in the Transitional zone. This improved boundary characterization is critical for clinical applications where precise delineation of overlapping anatomical zones affects diagnostic accuracy.
Results show that Rate-In’s information-theoretic dropout rates adjustment yields more accurate and clinically meaningful uncertainty estimates. Its ability to retain focused uncertainty while preserving anatomical detail makes it particularly valuable for medical imaging applications where precision and reliability are paramount.
Figure 4 illustrates the impact of dropout rates on information loss across network layers. In a pre-trained U-Net model, the decoder layers lose about 40% of information as dropout rates increase from 5% to 20%, while encoder layers remain more stable. This suggests that Rate-In effectively allocates dropout rates to minimize information loss where it is most critical.
Figure 4: Decoder layers are more sensitive to dropout-induced information loss than encoder layers, highlighting the need for layer-specific dropout rates. Analysis of information retention across U-Net layers as dropout rates increase. Decoder layers lose more information, while encoder layers remain more stable. This asymmetry supports Rate-In’s adaptive approach, assigning lower dropout rates to sensitive decoder layers and higher rates in robust encoder layers for efficient uncertainty estimation.
4.3 Out-of-Distribution (OOD) Analysis
Corrupted test images were generated following the setup and evaluation protocol of [9]. Robustness was assessed per corruption type and overall by averaging performance across types and severity levels. Rate-In improves robustness on MedMNIST-C, maintaining higher accuracy and AUARC across all OOD corruption types compared to the benchmark (Table 3). This improvement is especially notable at higher dropout rates and information loss thresholds, where benchmark accuracy declines more sharply. Rate-In mitigates these drops, demonstrating greater stability across datasets. A detailed breakdown by dropout method and corruption type is in Table 5 (Appendix B).
Table 3: MedMNIST-C Robustness Results: ACC(AUARC)
Dataset Method Full-Model $p\text{,}\epsilon{\text{=}}0.05$ $p\text{,}\epsilon{\text{=}}0.10$ $p\text{,}\epsilon{\text{=}}0.20$
TissueMNIST-C Rate-In .65 .63 (.77) .59 (.71) .45 (.48)
Benchmark .51 (.60) .34 (.35) .12 (.07)
PathMNIST-C Rate-In .54 .52 (.62) .50 (.59) .39 (.49)
Benchmark .49 (.59) .39 (.49) .25 (.31)
BloodMNIST-C Rate-In .77 .77 (.87) .75 (.86) .66 (.79)
Benchmark .77 (.87) .73 (.84) .59 (.73)
4.3.1 Computational Complexity Analysis
Rate-In’s dropout rate optimization has complexity $O(N_{\max}\cdot(n+f))$, where $n$ is the number of input instances, and $f$ is the complexity of computing information loss. Each iteration involves $O(n)$ dropout operations and $O(f)$ calculations, with dropout complexity further decomposing to $O(n\cdot d)$, where $d$ is the number of elements per instance. After optimization, inference runs at standard MC-dropout speed. See Appendix B for details.
Empirical results show that classification tasks (2D data, smaller models) complete within 0.70 seconds per instance, while segmentation tasks (3D data, larger models) take a few seconds to 90 seconds (Appendix B, Tables 68).